OpenStudy (anonymous):

Can someone please explain substitution to me? I understand how it works, but could you maybe give me a fast, efficient, way of doing it without just wasting all your time guessing random numbers that won't work...? Example: 6x+4y=2,000 x+y=410 Thanks so much! I really appreciate the help! :)

jimthompson5910 (jim_thompson5910):

substitution means replace one thing with another in math, it means that you replace one expression with an equivalent expression or number for example, if you are given x+2 = 5 then we can solve for x to get x = 3 to check this answer, we `substitute` 3 in for x. Ie, we replace x with 3 x+2 = 5 3+2 = 5 ... replaced x with 3 5 = 5 ... equation is true, solution x = 3 is confirmed

OpenStudy (caution0):

substitution is A way to solve a linear system algebraically. The substitution method functions by substituting the one y-value with the other.

OpenStudy (anonymous):

But what if you had two different variables? Not just x but with y too?

jimthompson5910 (jim_thompson5910):

when it comes to systems of equations, we can use substitution to solve first pick on an equation that you can solve for x or y without involving fractions the equation `x+y=410` is a good choice, so let's solve this for y x+y=410 x+y-x=410-x y = 410-x now we `substitute` 410-x in for y of the first equation \[\Large 6x+4y=2,000\] \[\Large 6x+4\color{red}{y}=2,000\] \[\Large 6x+4(\color{red}{y})=2,000\] \[\Large 6x+4(\color{red}{410-x})=2,000\] notice in the last step we replaced all of y with 410-x from here, you would isolate x using the previous rules you learned in algebra

jimthompson5910 (jim_thompson5910):

the reason why we're able to replace all of y with 410-x is because earlier we found y = 410-x y and 410-x are equivalent, ie the same thing

OpenStudy (caution0):

jim seems like he got this since hes beating me to the questions

OpenStudy (caution0):

lml

OpenStudy (anonymous):

Haha caution0, but thanks to both of you guys! It kind of pisses me off that I'm learning more from the internet in 5 minutes than in an hour at school...

OpenStudy (caution0):

lmao no problem ^.^